OpenStudy (anonymous):

Please Help!!! Solve the following system of equations algebraically 3x - y = 0 5x + 2y = 22

OpenStudy (anonymous):

Does it matter if it's substitution or elimination?

OpenStudy (radar):

from the first we get y = 3x

OpenStudy (radar):

substitute for y in the 2nd equation. 5x + 6x = 22 11x = 22 solve for x and then solve for y.

OpenStudy (anonymous):

Its says : Multiply one of the equations by a constant in order to solve by elimination.

OpenStudy (radar):

Oh, you just said algebraically.

OpenStudy (anonymous):

I mean thats what the question says...

OpenStudy (radar):

Multiply the first equation by 2 then add and solve for x.

OpenStudy (anonymous):

Okay so you were right. 11x = 22 so x = 2

OpenStudy (radar):

6x - 2y = 0 5x + 3y =22 ------------ 11x = 22

OpenStudy (radar):

Yes very good @Tesxtoast

OpenStudy (anonymous):

you help me finish! thanks so much!

OpenStudy (radar):

You're welcome, good luck with your studies.